A bank accepts thirty percent of loan requests from farmers. Ten loan requests from framers are randomly selected for consideration by the bank. Find the probability exactly 2 are accepted and more than 3 are accepted.


1
Expert's answer
2021-05-23T16:39:45-0400

The probability that exactly 2 are accepted is:

"P(2) = C(10;2)\u00d70.3^2\u00d7(1 - 0.3)^8 = 10!\/(2!\u00d78!)\u00d70.09\u00d70.7^8 = 0.2335."

The probability that more than 3 are accepted is:

"P(>3) = 1 - P(0) - P(1) - P(2) - P(3) = 1 - 0.7^10 - 10\u00d70.3\u00d70.7^9 - 45\u00d70.3^2\u00d70.7^8 - 120\u00d70.3^3\u00d70.7^7 = 0.3504."